A BIT OF OLD BATHGATE

AND

SOME OF ITS WORTHIES

—————

By John Stirling

A Bit Of Old Bathgate.

—————

At the regular meeting of The Lodge of Torphichen-Kilwinning. No. 13, held in the Temple, Jarvey Street, Bathgate, on Tuesday, 15th February, 1927, Bro. John Stirling, Secretary, delivered a lecture entitled “A Bit of Old Bathgate: some of its Worthies.” There was a large attendance of the Brethren, presided over by Bro. T.K. Irvine, R.W.M. The lecturer dealt principally with Jarvey Street, Main Street, The Hill, Cochrane Street, and High Street, bringing under review a number of “old characters” who lived, moved, and had their being in these areas. Bro. Stirling said —

Right Worshipful Master, Wardens and Brethren,

In acceding to the repeated demands of the Right Worshipful Master to give a few reminiscences of Bathgate, I have at last with much reluctance complied. I have decided to entitle my subject — “A Bit of Old Bathgate: Some of its Worthies and Associations.” Anything of a historical or antiquarian nature is always considered “a dry affair,” but I shall endeavour to make what I have to say as bright and cheerful and as entertaining as possible.

At the outset, I may almost say that the place whereon we stand is holy ground, this Temple having been built on the site of the old Hopetoun Lodge, No. 181. My grandfather was the Worthy Tyler of this Lodge, and as a boy I had free access thereto. It was from his lips that I learned how “Auld Skyte for Skyte” acquired his nickname. There had been a quarrel in Jarvey Street, and the accused men were being tried in Linlithgow. The Sheriff requested the witness to state what he knew of the case, when the witness replied — “The tain gaed the tain a skyte, and the tither gaed the tither a skyte, and they skytted awa’ at yin anither till they skytted yin anither doon,” a reply which created great laughter.

One of my oldest recollections in connection with the old Lodge is when my aunt got married to Kenneth M^{c}Kenzie, a glass engraver in the old Bathgate Glass Works. He was considered an expert of his art, and I have reason to believe that the beautiful decanters which adorn the Master’s table on Harmony nights, and presented to the Auld 13 by Bro. William F. Davidson, P.M., were examples of his handicraft.

Where Cavalry was Housed.

I shall not attempt to give you a description of this beautiful Temple — a brief sketch of which appears in the Bye-laws of the Lodge — but will ask you, in imagination, to accompany me in a brief pilgrimage to the various places I intend to mention. We leave the Temple and proceed to the end of Jarvey Street, where, until recently, the National Bank had its Offices. This in my young days was occupied as a hotel. It was here where the cavalry regiments, when shifting from one barracks to another arrived when on the march. The officers were billeted in the hotel, and their horses were stabled behind, the troop being dispersed and billeted in different parts of the town. They “fell in” opposite the hotel next morning, and it was a great treat for the boys to see them depart. One of the finest sights I ever witnessed in this connection was when the Dragoons were coming in the Muir Road, on a bright summer afternoon, the sun shining on the brass helmets made them glitter like a sheet of gold, the mounted band playing a march in which was incorporated the tunes “Auld Robin Gray” and “My Love she’s but a Lassie yet.” It made one feel they’d like tae gang and be a sodger. We had visits at different times from the Inniskilling Dragoons, the Lancers, the Hussars, and Scots Greys — Bro. Thomas Shanks, of Auld 13, being a Sergeant in the latter.

We will now cross the street to The Old Curiosity Shop, presently occupied by a “far famed and well known” printer (Bro. Stirling) — the favourite haunt of a large number of members of the Craft, where many jokes are told and cracked in the “works” (back shop). This was at one time a chemist’s shop, occupied by Mr Stuart Brown, whose son quite recently purchased the estate of Cathlaw. He laid the foundation of his fortune in the wee shop in Jarvey Street. Methinks I hear someone say — There are still fortunes being made in that queer auld farrant place.

Next door, at one time, was housed the Volunteer Armoury. It was here where I took the “Queen’s Shilling,” some four decades ago. The drill instructor’s name was Nathaniel, his by-name being “Old Nat.” He was a member of the old school and one of the finest instructors in the Volunteer movement. I am pleased to say we have two of his grandchildren present to-night — Brothers Thomas and Peter Nathaniel.

The shop now occupied as an Italian store was the business premises of the late Bro. Alexander Davidson, plumber and tinsmith; he was the worthy Secretary of this Lodge for many years.

Proceeding along Jarvey Street, we come to the Corn Exchange. The Bathgate Market was held here, horses and carts lining the street from the Bunker (Brown’s Square) right along to Bryson’s Market Inn at Mid Street Corner. The cereals were sold in the Corn Exchange. It was here also that the farm servants were hired.

In what was known as the New Kirkyaird, adjacent to the present Parish Church, lies the body of Geordie Puddin’, the Bathgate gaberlunzie — a most notorious character. The gravestone was erected by J. Calder, builder, and paid for by Thomas Durham Weir of Boghead House.

The house presently occupied by Bro. R.T. Arbuckle, dentist, was the National Bank premises before being transferred to the other end of the street. John Johnston was the Bank Agent; he also carried on a licensed grocer’s shop here, and farmed the farm of Ballencrieff — “the east side of the burn.”

On the opposite side of the street, where the Cooperative Buildings now stand was the Hopetoun Inn and hiring establishment, which at a later date was occupied by Dr Corbett, who is credited with having called on John Newlands while on a visit to Jamaica and suggested to him to remember his native town in his will. Dr Corbett was a R.W.M. of Hopetoun Lodge, No. 181. The Little Advocate (Bro. James Gardner), resided here for a number of years.

Jack Robertson, the auld cobbler, lived next door in a two-storey thatched house, which was nearly set on fire at a Municipal Election, by a “paraffin ball” alighting on the roof. Luckily, the fire was put out before it got a firm hold.

In those days the voting took place in the Corn Exchange, the inhabitants turning out in hundreds to await the result. The victorious candidates were carried shoulder high to the nearest public house, where they had to do the “handsome.” The public houses did not close till 11 p.m.

Turning into Main Street, we pass what was the residence of auld Jock Murray, nick-named the Polar Bear. It was a three-storey tenement, known as the Blue Tower, deriving its name from the fact that it was the first house in Bathgate to be roofed with blue slates. Passing up Market Street we arrive at the Hill, where stood the Jail, in which was accommodated from time to time such worthies as Singing Jim, the Rolling Eye, Chic Chic, Katie Fuff, Dumbarton Jock, the Whale, Hen Jock Bird, and Geordie Puddin’.

Later Characters in the locality were Bobby Blue, Pipes o’ Rory, Tidy Fa-lal, Sally Sharky, Nanny Douce, Rab Tickler, Stulty Donal’, the Pitcher, the Weasel, the Hoolet, the Rat, the Stoup, Bobkin, Molly Greens, the Jingler, Gad Wull, and Susy Dunbar.

Jock Bird and Geordie Puddin’ were locked up on one occasion at the same time, having quarrelled as to who was to be boss at the digging of a drain.

They hadna been lang at their wark,

No o’er an hour or twa, Till there arose as wild a shine As mortal ever saw.

Wha wad be maister coost them oot, Which spoil’d a’ wark thegither;

As Puddin’ flung a sholfu’ up, The Bird flung doon anither. Jock made a spring at Puddin’s throat

Wi’ firm, determined grip, And savage, savage was the fecht Till Pillans locked them up. An’ a’ that day, an’ a’ that nicht, The Hill was kept in steer, For when they werena singin’ psalms

They baith fell tae sweer.

The Jail was latterly occupied as a storage for the scavengers’ besoms and clauts. It is now one of those open spaces which Bro. Irvine is so anxious to get established throughout the town, swings having been erected for the children.

Alexander Marjoribanks.

To the right, some three or four hundred yards distant, stands the beautiful ancestral home of the Marjoribanks’, designed by the Bros. Adams, famous world wide known architects. Early in the 18th century the estate of Bathgate passed from the House of Hopetoun to that of Marjoribanks. To Alexander Marjoribanks the people of Bathgate owe a deep debt of gratitude. He insisted on contesting the will of John Newlands against his heirs, who wanted the will set aside, and he gave a guarantee that he would bear the expense of the litigation. The case was fought in the Court of Session, and decided in favour of the town. It is said that the amount of money left was over £60,000, but after expenses had been paid there was only £15,000, which went towards the building of the present Academy, the £15,000 being only the interest on the principal for ten years.

Mr Marjoribanks was a very generous gentleman, was known as “the good old laird,” was a good friend to Bathgate, was elected the first Provost of the town, and deserves to be ever held in grateful memory. His grandson is at present the minister of Stenton, Prestonkirk parish, who is in possession of his grandfather’s pruning knife, a relic of bygone days. The family are also in possession of numerous oil paintings and Biblical engravings which at one time adorned the walls of Balbardie House. It is a great pity that the estate passed out of possession of the Marjoribanks.

John Newlands.

About twenty yards to the right of the old Jail is the site of the house where was born John Newlands, founder of Bathgate Academy. The history of his going to Jamaica, making a fortune, leaving same for the education of the people of Bathgate is well known. He was born on 17th April, 1737 — that is the recognised date, although there has always been some doubt about it — in fact, it has now been proved that that was the baptismal date. The house was long occupied as a joiner’s shop by John Boag.

Sir James Young Simpson.

About the same distance to the left from where we stand is the birthplace of Sir James Young Simpson, the discoverer of chloroform, who was born on 17th June 1811. His father was a baker, and his mother a good pious woman, endowed with a vast amount of common sense. One day (as a boy) when the future Professor came into the house, with a big hole in the heel of his stocking, she took him on her knee and darned the stocking, remarking — “My Jamie, when your mother’s awa’, you’ll remember that she was a gran’ darner.” At school he was of a steerin’ disposition, and was possessed of a remarkable memory; before and after school hours he had to go to the baker’s “brod” dispensing “baps” and scones to his father’s customers. I will not trace his university career, which was paid for by his sister and brothers clubbing together. In 1832 he graduated as M.D., and was appointed to the Chair of Midwifery in 1840.

The great discovery of chloroform was made in the year 1847, and was hailed with acclamation and enthusiasm throughout the world. He died at Queen Street, Edinburgh, on 6th May, 1870, in his 59th year, and never was man more lamented by all ranks and classes of society. He was buried on the southern slopes of Warriston Cemetery, the spectators being estimated at over 100,000. His funeral was a great and solemn ovation. The family were offered a burial place in Westminster, but they wisely decided to bury him in the city where he laboured with so much acceptance. If ever man was a friend of the poor — that man was Sir James Young Simpson.

I have mentioned the names of three gentlemen — in fact, I may say three great philanthropists — two of whom done a great deal for Bathgate, and one of them inestimable good to mankind; still, Bathgate has not yet seen its way to erect a memorial in memory of any of them. So passeth away all earthly glory.

Proceeding down Main Street we pass the Tontine Close, and the once famous Collogie — in which was a place where the “three balls” held sway — you all know what that means. It was here that Wallace made his “ha’penny dips,” the workers being known by the nick-name of “cracklins.” We also pass three public houses in close proximity — some twenty yards separating them. In one of them, occupied by Mey Fyfe, auld Gad Wull and his son were in “killin’ their craw.” The son ordered a quart of half-and-half. Auld Gad asked — “What’s that ye’ve ca’d, George?” “Beer, faither,” said George. “O man, ye shouldna ca’d that; there’s nae maut in’t,” quoth the father. Charlie Morrison’s bake-house, shop and dwelling-house stood at the right hand bottom of Main Street.

Dragoons and the Minister.

At the foot of the hill stood the Auld Kirk. A great dispute took place here at the ordination of a minister, in 1717, against the wishes of the people. This is best told in the following extract :–

“The ministers serving the edict, knowing the badness of their cause, and the evil part they were acting, thought not fit to do it until they got a troop of Dragoons to be a guard to them, and accordingly on 17th November 1717, being approaching the town, they caused beat their drums, and draw their swords, and in this position came through the town, guarding the ministers into the church, riding and striking with their naked swords, at the women and others standing gazing upon the wayside, which was a melancholy Sabbath in Bathgate, the Sabbath day being much profaned.”

The Kirk stood then doon at the Hill,

And stands unto this day; For lang it was a whisky shop,

And roarin’ trade did dae. An’ awfu’ rumpus here occurred

At placin’ of a curate, Which fired their Presbyterian bluid,

And made them quite infuriate. They took their cue frae Giles’s Kirk,

And famous Jenny Geddes; The curate, fley’d, ran up the Style,

Pursued by wives and laddies. A captain wi’ a troop of horse

Was posted no far by, Wha gave command tae charge the crowd

That on the Hill did lie In the melee a horseman fell,

And ither twa were wounded; While maister curate fled the toon,

The cure ower hot he found it.

The church property was long in the possession of Mr Hume Chalmers. A cut from the sabre of one of the Dragoons was long to be seen upon the door. Tradition points out a tombstone within the old ruin of the Auld Kirkyaird, bearing a large sword, said to be the tombstone of the Dragoon who was killed in the melee. The Guard House, in which some of the Dragoons were stationed, still stands at the corner of Dykehead Lane, and is known as the Guard House to this day. Happily we now live in much different and happier times.

Professor Diney.

We shall now descend Cochrane Street, passing Whippergigg Wynd and Kamehead, till we come to the house of Sandy Christie, one of Bathgate’s most famous curlers, who kept a licensed grocer’s shop. He was succeeded by John Forrest (“Jock Purr”), who for the long period of close on thirty years acted as Treasurer to the Bathgate Academy Procession. In this property resided Alexander Hamilton, a famous barber — alias Professor Diney. The Professor, in addition to being a master in the tonsorial art, was a keen draughts player, and it was difficult to snatch a game from him.

The late Bro. Dr Kirk related the following story to me :– Wyllie, the “Herd Laddie,” champion draughts player of the world, was on a visit to Bathgate, and residing with the late Bro. Dr Kirk’s father. The old Doctor had a visit to make at Avonbridge, and on the way he called at Diney’s and left the Herd Laddie till his return. After a short conference, Diney asked the Herd Laddie if he would like a game at the draughts. The champion said he would; he had tried his hand at the game but was not much of a player. Diney won the first three games, and was in ecstasies. The Herd Laddie won the next six games in succession, when Diney, exasperated, jumped from his seat and declared that “he was either the Herd Laddie or the deil himsel’.” The Professor shaved for a penny, and had the following names for his razors :– Meadow Queen, Scotland Yet, Honey Bee, Rattlesnake, and the Rasper.

After all, there was nothing wonderful in Diney shaving for a penny. I myself have had many shaves for a penny by Jock Newlands, who resided in Hopetoun Street. I remember on one occasion, my wife having had a long illness, Jock, out of stark love and kindness, having shaved me I proffered the customary penny he said – “Nevell mind, John, ye’ve had a lot tae dae the noo; keep ye’ll penny.” The following verses are by Poet Shanks :–

Half way betwixt the Prison Hill,

And what was yince the Bathgate Mill, There lived – nay, there is living still, The great Professor Diney. A spotless apron, pure and white, Hangs gaily owre his gurdy kyte, In sark sleeves, morning, noon and night, Goes great Professor Diney. Wi’ blandest smile and easy grace, He tak’s the handle o’ your face; The king o’ a’ the barber race

Is great Professor Diney. Syne what a loving way he’s got O’ lingering about ones throat; Ye Gods! if drunk, or mad, what not, Might do Professor Diney. For stylish cut, and quick dispatch, In Scotland braid there’s not his match; And who a game at draughts can snatch From great Professor Diney.

We will now retrace our steps up Cochrane Street and Main Street, till we reach the Cross Keys public house, now occupied by Mr Webster as a confectioner’s shop. Next door to this was the baker’s shop of Bailie Alexander Russell; it is now occupied as a private dwelling. Bailie Russell served in the Town Council for the period of 35 years. We now reach :–

The Sun Inn.

On the 25th January 1892, there passed away in the old “Railway Inn,” James Bowie, who for 90 years, had passed his days and nights under its roof. The house belonged to his father, and in his time was called “The Sun Inn,” a flaming representation of which was placed over the door, with the legend or motto underneath — “The best whisky under the sun.” It must have been good and pure, for Dr Kirk (the late Bro. Kirk’s father) always recommended his patients, when prescribing spirits, to go to James Bowie’s. He was rarely called by his baptismal name, but as “Provost Bowie.”

Provost Before he was Born.

The origin of the title given to him was thus :– Previous to the erection of Bathgate into a Burgh, Tom Dick, afterwards Town Clerk, and the Laird of Marjoribanks and Balbardie (of sainted memory) were in the “Sun Inn” discussing the prospects of securing from Parliament the much coveted Act. Mr Marjoribanks said, “I think, Mr Dick, if we secure the Act from Parliament, that we might, with great propriety, make the next son born to Mr Bowie our first Provost.” “Agreed, Laird, agreed,” said Tom Dick. When a son was born to the host of the “Sun Inn,” Mr Dick hastened over to “the big hoose” and requested the Laird to come and see Bathgate’s first Provost.

That was the origin of the title, and Bowie used to say, “I was made a Provost before I was born.” In addition to carrying on the Inn, he wrought fields near the farm of Hardhill. He always wore a satin or “lum” hat, and visited the fields in that garb.

A lover of dumb animals, he had a number of cats and dogs; he also had in his possession a monkey which was a gifted character. It was known as “Batty Bowie’s Puggy,” and it is said that the monkey used to get hold of one of the cats, plunge its paws into hot soup, in search of a bone for his satanic majesty — a sure way of preserving his own skin. The puggy as it grew older showed signs of a ferocious nature, and was presented to the Edinburgh Zoological Gardens, much to the relief of the maids, cats and dogs, at whom he was continually snapping. Years after when Bowie paid a visit to the Gardens he observed a cage with a warning notice for visitors to keep back, as the monkey was dangerous. Notwithstanding the warning of the keeper the Provost entered the cage, when the monkey recognised his old master and went wild with joy. Amongst his other favourites was a long-legged black and white “soo,” and a “wee timorous beastie” which he fed daily. His horses were named Rachin, Dobbie, Haddington Horse, and the Priest.

Bowie was a keen Freemason, and was a member of “Auld 13” for the long period of 63 years. The toddy tumblers at that time were rather narrow at the bottom, and were frequently capsized. At the instrumentation of Bro. Bowie several dozens much broader bottomed tumblers were secured — the only remaining one left in 1892 was in possession of the proprietor of “Ye Howff,” Bro. James Wallace, a small chip off the edge being the only damage it had sustained during the wonderful occasions when the Brethren were called from “Labour to Refreshment.”

The “Provost” was possessed of remarkable eyesight, and one of his accomplishments was that he could write the Lord’s Prayer on a space which a sixpenny piece would cover, and so good was his eyesight and so steady his hand that he performed this feat a few months before his death.

A Relic of Prince Charlie.

The “Provost’s” mother was the possessor of a precious relic of Prince Charlie, in the shape of a set of Royal Stuart Tartan Curtains, which adorned the bed on which the Prince slept, on the night in which he passed in the old Deans mansion. They were purchased by Mrs Bowie at the plenishing sale of the old mansion house, and they adorned two beds in an upstairs room of the Railway Inn for many years. One set was mutilated by parties cutting off portions as mementoes of the unfortunate Prince; £5 was offered for the other set, but refused.

One of the retainers of the Cross Keys and Railway Inn was Jock Shirra. For over forty years Jock assisted in dispensing pies and drams — a special treat on the Fair days; latterly, when the tall and buirdly frame of Jock began to bend, he told his old employer that he was not “soople eneugh,” but the “Provost” and his customers had been so long accustomed to the cheery and good-humoured countenance of Jock that he was urged to “come aboot the hoose for luck’s sake,” and sit by the “fire en’;” and so to the end he was always a welcome visitor at the auld thack hoose.

The Bunker.

On entering Brown’s Square (the Bunker), we find Mary Cherry, Bathgate’s first and only lady carter; Mary was a hard working woman, and carted coals from Balbardie Mine to her various customers throughout the town. There also resided here Bee Rab, San Ponder and his cuddy, Jamie Kirkland and his cuddy, and “Coachy” Alexander, ostler to Batty Bowie.

The Bunker was once the gathering ground of the travelling shows — “penny gaffs” and shooting saloons, and was a favourite resort of the travelling acrobat. I have seen Old Malabar, a famous Glasgow character, perform here many times. The principal feature of his show was to throw a cricket ball as far up in the air as he could and catch it in a cup attached to his forehead, a feat in which he never failed.

Delaney, “Orr’s long clown,” also had his outside show here. His principal feat was to stand on a horse’s bare back — the horse galloping round the ring — and throw off waistcoat after waistcoat; it was generally considered that he had over two dozen waistcoats on. He died in Broxburn, and it is good to relate that he had a good friend there in the person of Bro. Norman Henderson, who was a native of Bathgate, and managing director of Broxburn Oil Company from its inception till his death.

That’s Him — That’s Starkie.

We shall now ascend High Street, known in the olden days as “Shuttle Row,” passing on the way the abode of Dr Kirk; the house occupied by Dominie Macgregor, who taught a day school at “Kallifat” and a night school in the old Masonic Lodge, Gideon Street; also the houses occupied by Thomas Dodds, solicitor, Dr Dickson and Dr Longmuir — now occupied by Bro. James A. Pow, the respected treasurer of this Lodge. We arrive opposite the abode of “Starkie,” the next house being occupied by the “Apostle” Fleming.

As up high Bathgate street ye spiel, An’ fore a wee snug theekit biel, You spy a queer auld farrant chiel, Stript tae the sarkie,

An’ borin’ pump wi’ cautious skeel — That’s him – that’s Starkie

John Stark, the far-famed and well-known “Starkie,” passed peacefully away ’mid the gloom of a December morning in the year 1882. Possessed of a kindly, cheery and obliging disposition, he was always ready to lend a helping hand to any scheme intended for the benefit or welfare of neighbours. He was never so thoroughly in his element as when sitting at the head of his plain, yet substantially loaded table, dispensing haggis and beef and greens, interspersed with “willie wauchts” from his stoneware bottle. He was a first-class maker of peeries, boys coming from all parts of the town for one of Starkie’s “wummers.”

A great and intense admirer of our National Bard, he always celebrated, along with a few kindred spirits, the Poet’s natal day, when he left pumps and peeries, wash tubs and spinning wheels, to take care of themselves. “The Immortal Memory” was always proposed by Mine Host, and never in any gathering was it drank with more enthusiasm. As the night wore on, assisted by John Barleycorn, they sung lustily “The cock may craw, the day may daw,” and “Auld Lang Syne” time about. I am creditably informed that the king of the song was auld Starkie, he being the last to fa’ beneath the table. Tom Anderson, the joiner, was always present at the Burns splores, and the event is immortalized thus by the Blind Poet of the Deans —

When Joiner Tam and he foregather, A kindred spirit — sic anither — They’ll sit far on for days thegither, O’er glass hobnobbin’; Their tongues it wad be vain to tether When loosed on Robin.

These twin enthusiasts agree That Robin fairly taps the tree; A’ ither bards maun bow the knee When Robin sings; He is, and ever more shall be, The King o’ Kings.

A great Academy Procession enthusiast, his house (which was a thatched one) was always decked on Procession day with nick-nacks of all conceivable kinds, among them being a model of “Burns at the Plough,” and “Jenny at her Spinning Wheel,” both of which were made by Starkie. He had a piece of calico about 3ft. square, which he never failed to hang out, with the verse printed thereon from Burns’ Address to the Brethren of St. James’ Lodge, Tarbolton —

A last request permit me here, When yearly ye’ assemble a’; One round – I ask it wi’ a tear —

Tae him, the Bard that’s far awa’.

The two following verses are by Bro. Robert Fleming —

Oor Burns enthusiasts will greet, When they reca’ the oors sae sweet, That they did spend when a’ did meet,

Tae weet their craggies, And fill their wames at Starkie’s fete

Wi’ famous haggis. His guns and pistols, jugs and skulls, His dirks and swurds, an’ auld snuff mulls, His picture books an’ ancient quills,

His nick-nacks a’, May a’ be scattered tae the hills

Sin’ Stark’s awa’.

I conclude my remarks on Starkie by quoting the undernoted verses from “The Auld Bathgate Worthies” —

Noo gane is auld Starkie, His nick-nacks and larkie, Tae whiten and moulder

In cauld, wormy clay; The bairnies, a’ bleerie, Nae mair dose their peerie, The auld Bathgate worthies

Are a’ wede away. Farewell for ever, Starkie! Perhaps thy wee bit larkie To thee in ither realms

Pours out its lay; Thy link has burst at last That bound us to the past — The auld Bathgate worthies

Are a’ wede away.

I do not intend to go any further with you at this time, but will take you down to the corner of Gideon Street, and leave you in “Ye Howff,” presently occupied by Bro. Tom Wallace — the meeting place of the once famous Breeches Club: a subject for a paper by itself. The only remaining member of that one time popular and select Literary Society is Bro. Robert Waddell, the worthy Bible-Bearer of Auld Thirteen. Let me here introduce you to an old Bathgate worthy —

James Thornton, the Drummer.

In June 1892, the grave closed over one of Bathgate’s well-known characters — James Thornton, better known as “Pousless.” Jims was a wiry wee body, and active. He was all out for peace, and woe betide the boys he caught turning on taps at wells, running away the water; he generally cuffed their ears and afterwards said “It’s a pelfeck disglace that a pack o’ scoondlels should be allowed to pack the tholofale,” and told them that he would hand them owre by, viz, the Police Office. He was a regular attender at the Burgh Court, and nothing pleased him better than when a severe sentence was imposed. Jims, for well nigh half a century, with his “dlum” ushered in the Procession morning playing “Fire in the mountains, rin boys rin,” or “The British Grenadiers,” his two favourite tunes. It was a common thing for boys in those days to be up on Procession Day any time between three and five o’clock in the morning. “Pousless’ was in the habit of attending the slaughter house. One day he got a pudding filled with blood, and rolled it round his neck. He went home and asked his mother for a penny, stating that if she refused he would cut his “thloat,” which he did, falling on the floor with the blood streaming from the pudding, his mother rushing out to the street exclaiming that “Jims had cut his throat.”

The following stanzas by Poet Fleming, a member of Auld Thirteen, depict Jims to a nicety :—

Wha’s yon wee bit bodie that steers up an’ doon, As if he was laird o’ the hale country roun’, Aye lauching’ an’ nodding, or hummin’ a tune? — Yon’s Jims – Jamie Tholnton, the “Dlummel.”

At meetin’ or market ye’ll fin’ Jamie there, At Coort or at concert, at kirk or at fair; Nane move a mair consequential air

Than wee Jamie Tholnton, the Dlummel.”

Like the gallant “John Murray,” he’s well versed in law, An’ wi’ Jims it was whyles unco dangerous to thraw; Gin his heid begoud shakin’, ’twas time to g’awa’,

Or ye sune cam’ to grief wi’ the “Dlummel.”

But noo since his auld frien’ the “Shillah” is deid, Jims disna sae muckle wi’ the law fash his heid; The Coort-room to him is a cauld place indeed,

For nae Lordship shakes hands wi’ the “Dlummel.”

As “Protector of Peace” there didna reside, When young, Jamie’s equal in this country side; Oor polismen noo are sae slow i’ the “stride,”

They never could cope wi’ the “Dlummel.”

If a laddie had touched a well on the street, An’ Jims chanced to be on his every day “beat,” It was “charge,” an’ a rin boys rin “retreat,”

Or a “cuff ’o the lug” frae the “Dlummel.”

In his palmiest days, oor “loafers” fared bad, For a stamp wi’ his fit gar’d a’ quickly “pad” — A “pelfect disglace” that men “dlunk an’ ill clad”

Should lounge on oor streets, quo’ the “Dlummel.”

An’ as for a “baker” auld “Charlie” could tell, There wisna a baker wi’ Jims could excel; For carryin’ heidfu’s, piled up by the ell,

“Lobie White” had nae chance wi’ the “Dlummel.”

An’ as the toon’s “dlummel,” he hisna a peer, Frae the Cannibal Isles to the toon o’ the “Queer;” His notes are sae true, aye, sae sweet, and sae clear —

A “model” musician’s oor “Dlummel.”

Folk bounce and blaw aboot yin Johnnie Bain, While some praise Jock Gardner, or Alec M^{c}Lean — It’s bombast! There’s nane here, or e’en ower the main,

Could “play the ae side” o’ oor “Dlummel.”

James Forrest.

I cannot close without making a short reference to one who was much loved and respected by all who know him — the late Bro. James Forrest, the much esteemed Bard of Lodge No. 13. To his flawless integrity in all relations of life, to his geniality and generosity in every sense of the word, and to his wholly delightful urbanity and charm of manner, all who knew him bear unfaltering testimony.

James Forrest was no ordinary man. No one in the district had a better knowledge of the folk lore and ballad literature of Scotland; he had a keen appreciation of Scottish humour, and had a never failing budget of anecdote. He had a most extensive knowledge of old Bathgate and its worthies; it was a special delight to sit and listen to him relate stories of old Bathgate scenes of long ago. He was endowed with a happy temperament, a never failing fund of jollity and good humour, and possessed of fine literary taste. One of the finest traits in his character was a love for dumb animals; he was well versed in the habits and haunts of the birds of the district. He was tall in stature, and in every respect of the word a gentleman.

His life was gentle; and the elements

So mixed in him that Nature might stand up And say to all the world: This was a Man.

Bro. Forrest was the Secretary of the Beeches Club from the time of its inauguration till its decease; his reading of the minutes was a literary treat, and often much more enjoyed than the essay or lecture.

Possessed of much literary ability, he was the author of “The White Lady of Caputha’ Bogs,” a short story which appeared in the local press; “Bird Notes,” “Scotland’s Ploughman Bard,” “The Puir Wife’s Brae,” “A Day Among the Haws;” “The Old Firm at the Old Address,” being a graphic description he paid to the old Livingston Inn, accompanied by the Poet of the Deans; “The Death and Burial of the Thrush;” and a poetical effusion dedicated to his Mother Lodge, Auld 13, sung to the tune of “Scotland Yet,” which I reproduce in full :—

Beneath Kilwinning’s auld roof tree,

Assembled are we a’, To spend the hours in social glee,

And drive dull care awa’; For round about this board to-night

True Masons all are we — Dear Auld Thirteen, dear Auld Thirteen,

Our Mother Lodge is she; May fortune smile upon her Sons

Wherever they may be.

Within this mystic Temple, then,

Let faith and hope entwine; May Charity’s effulgent robe

Clothe all in love divine, And Wisdom wait upon the Sons

Of Light and Harmony — Dear Auld Thirteen, dear Auld Thirteen,

Our Mother Lodge is she; May fortune smile upon her Sons

Wherever they may be.

Amid the changing scene of life,

’Mid worldly strife and care, Let all our actions fashioned be By the Compass and the Square: And, mindful of Life’s Golden Rule,

Join Love and Unity — Dear Auld Thirteen, dear Auld Thirteen,

Our Mother Lodge is she; May fortune smile upon her Sons

Wherever they may be.

Of orders kings and nobles boast,

Of starts and royal blood, Antiquity is stamped on ours,

It dates from near the flood; O’er all the world our Order’s known

By deeds of Charity — Dear Auld Thirteen, dear Auld Thirteen,

Our Mother Lodge is she; May fortune smile upon her Sons

Wherever they may be.

Then to the Craft let’s pledge a toast, With honours three time three, Auld Thirteen’s Sons in every land, True Masons may they be:

They’ll find across Life’s troublous scene

A blessed eternity — Dear Auld Thirteen, dear Auld Thirteen,

Our Mother Lodge is she; May fortune smile upon her Sons Wherever they may be.

He was laid to rest in the Auld Kirkyaird at Kirkton, on Friday, 7th July 1916, a deputation from the Lodge of Auld Thirteen paying him the tribute of following his remains to the last scene of all — that borne from which there is no returning. I have no hesitation in saying, and I conscientiously believe, that James Forrest would face the Grand Architect of the Universe with a clear conscience, and that he would receive the “Well done, good and faithful servant.” “After life’s fever, he sleeps well.” In the words of A.M. Bisset —

Nae mair he’ll spiel the Puir Wife’s Brae At dewy dawn or gloamin’ grey, To tune his heart to Doric lay, An’ muse alane;

For doon the gait we a’ maun gae Has Forrest gane.

A great many of Bathgate’s old actors have made their exit from life’s stage: the lights have been turned off, and the curtain rung down.

Vale! old friends, take you for all in all, when shall we look upon your like again?

A Relic Of The ’45.

—————

ADDENDUM

To the Lecture on “A Bit of Old Bathgate.”

—————

At the regular Meeting of The Lodge of Torphichen-Kil-winning, No. 13, held in the Temple, on Tuesday, 3rd May 1927 — Bro. T.K. Irvine, R.W.M., presiding — Bro. John Stirling said :— Right Worshipful Master, Wardens and Brethren, — In the recent address which I gave on Bathgate Characters, I made reference to the curtains which were in possession of the late Bro. “Provost” Bowie, and which adorned the bed on which Prince Charlie slept in the Deans Mansion House, or more correctly speaking, Boghall House. The mansion was the property of Colonel Norvell, and was occupied by the Norvell family. I lost all trace of the curtains after the death of the “Provost,” and at one time thought they had gone into possession of John Macnab of the Glen, he having procured the old hostel occupied by Bowie.

I am pleased to say that I am now the happy possessor of those historical relics, Mrs Kirk having very kindly presented them to me. They came into possession of the late Bro. Dr Kirk at the death of the “Provost.” They are made of the best worsted material, of the Royal Stuart tartan, and although close on 200 years old, are in a fairly good state of preservation. The visit of the Prince to the locality and the history of the curtains are mentioned in “Rambles Round Bathgate,” by Mr Convery, and are chronicled in the “Rhyming History of Bathgate,” by Alexander Hamilton, the poet of Kirkroads, as under :—

This brings us tae the ’Forty five,

’Tae Charlie and the Clans, Wha slept ae nicht on Bathgate Hills

When gaun tae Prestonpans. The rank and file, row’d in their plaids,

Lay doon at Clinkinstane, While Lord George Murray wi’ the Prince

Unto the Deans has gane. A royal banquet there was spread,

Wi’ Norvell at its head, Wha drank tae Johhny Cope’s defeat

Before they gaed to bed. A worthy Provost in our toon,

Within his house can show The curtains o’ the Prince’s bed,

The counterpane an’ a’. At dawn o’ day the Cameron Clan

Brak’ in on Lizzie Meikle, And toom’d her girnal and her kirn,

Which put her in a pickle. Some ran wi’ jugfu’s o’ the cream,

And made it into crowdie; While others clap’d theirs on the fire

For brochin and powsowdie. Brave Lizzie ran straight to Lochiel,

And telt him ’bout his men; The Chieftan laugh’d, and frae his purse

Drew gouden pieces ten, And flung them right intae her lap,

Then turn’d and wish’d good-bye; And lang did Lizzie Meikle tell

She ne’er made mair aff kye.

The Deans had in those days a fine baronial hall. It formed part of the ancient Barony of Boghall, and at that time there were between thirty and forty crofters on the estate; there was also a smithy, which was famed for the manufacture of tackets. There was also a thirling mill, and the course of the old mill lade could be traced up to quite recently.

Norvell, who presided at the Prince Charlie banquet, was Laird of the Deans; the worthy Provost referred to, and who could show the curtains of the Prince’s bed, was Bro. James Bowie, of the Railway Inn; Lizzie Meikle, whose “girnal” the Cameron Clan plundered, was the great grandmother of Alexander Hamilton; crowdie, brochin, and powsowdie, are mixtures of milk, meal, and boiling water — forming a kind of drummoch or “het-gab.” It is well known that the Prince’s army, when on the march, stole goods and food wherever they could; and when found fault with by the people whom they robbed, the soldiers said they would be recompensed when “Charlie cam’ in tae his ain.”

Prince Charlie.

Prince Charles Edward Stuart was born at Rome on 31st December 1720, and was the grandson of James VII. of Scotland. At the age of 22 he conceived the idea of recovering the throne of his ancestors. He landed from France in the Highlands of Scotland in 1745 with seven attendants, and was joined by several chiefs and their vassals. The force of Charlie was small considering the great work that he was attempting, but he was not deterred from advancing.

He left Perth on 10th September of that year, his followers rapidly increasing, and crossed the Forth above Stirling. He passed the night of the 15th at Callander House; passed Linlithgow Bridge on the 16th, and proceeded through Linlithgow and Kirkliston towards Edinburgh. It was while on this journey that the Prince slept in the Deans Mansion House. He captured Edinburgh, and afterwards gained the battle of Prestonpans. He penetrated into England. His rapid advance made the British Government tremble. In not making straight for London he lost the “move.” He decided to retreat when at Derby, and arrived in Glasgow at Christmas.

We shall always admit the loyalty displayed by Prince Charlie’s followers in his unfortunate advance. If he had pushed on for London, after the panic caused by his first success, instead of wasting precious time in Edinburgh, there is no saying what might have happened. The weeks he spent there lost him the Crown. When he did decide to advance, it was too late; for his enemies had got over their panic, and were fully prepared for resistance. Even at Derby if he had persevered he might have been successful. Although all hope was lost on Culloden Moor, there is a charm about the adventure which cannot be extinguished. The sad and pathetic ending of his career is well known history, and need not be recorded here.

Upon invitation the Prince slept one night, if not two, at the Mansion House of Boghall at the Deans, at that time the seat of the Norvell family.

Portrait of Prince Charlie.

Mrs Kirk is in possession of a beautiful oil painting of Bonnie Prince Charlie — considered to be an excellent likeness — which also came out of Boghall House, and was the property of the Norvell family. It was purchased at a displenish sale — at which the curtains were also sold — by Mr Joseph Pearson, who afterwards gave it to a gentleman in Edinburgh named Stuart, who was supposed to be closely connected with that Royal House. The Pearsons, who removed to Australia, were related to the Shanks family.

The picture ultimately found its way back to the Deans, when it came into possession of the family of Poet Shanks. When the poet removed to Kirkton Lodge it occupied a prominent place there. It is supposed to be an authentic portrait of the unfortunate prince. The Shanks family firmly believed it to be so, and coming as it did from the Norvells, who had strong Jacobite tendencies, it is more than likely to be the case.

I was told by the Poet that the picture was to be left to Dr Kirk, the conditions being that the Doctor was to attend the Poet professionally till his death — a duty which the Doctor nobly fulfilled. When the Doctor received the picture he had it renovated and reframed. I saw and admired it many times at Kirkton Lodge — the Poet never forgetting to tell me where its resting place was to be — and I had the pleasure recently of again seeing it, through the kindness of Mrs Kirk, of Rowan Bank, where it is hung in the hall staircase.

I have heard it insinuated that the portrait is that of Charles II. It may be true. Not having met any of the gentlemen, I am not in a position personally to say which is correct. However, that is a question that could be easily settled by any inquisitive person.

————

The curtains were on exhibition, and all present showed keen interest in the historical relics.

Kirkton Mansion House.

It may not be amiss to give a short description of Kirkton Mansion House, the Lodge of which was so long the residence of the Blind Poet of the Deans. The Lodge was erected in the year 1844, and was demolished in 1927 by the Town Council of Bathgate — a new building being erected for the keeper of the Public Park. The Mansion House stood at the east end of Bathgate Public Park, near to the farm steading now occupied by Mr Young, and was of considerable magnitude, but was not so commodious as it looked owing to the thickness of its walls. It was two stories in height, with partially underground kitchen and cellars. There were several steps of stairs to the main entrance, and it had two wings at each end in which were several flights of stairs giving access to the upper apartments. There was no landing places at the top of the stairs, so that the entrance into the rooms was sudden and abrupt.

The building was demolished by the instructions of Captain Hart, the then proprietor, about the year 1862. Bro. Thomas Johnston, an old and honourable member of this Lodge, had the contract of taking down the mansion and removing the stones and material and disposing of them at his pleasure.

A tablet stone which stood over the entrance had inscribed on it the names “Thomas Sherp and Marion Dalmæhoy, 1599.” It is conjectured that these were the names of the then proprietor and his wife. The tablet stone is built into the gable end of an outhouse at Petershill Cottage, then occupied by Bro. Johnston, where it still remains, and can be seen by any person passing, the gable facing the road. There is also above the tablet the figure of a cherub which was brought from the mansion.

There was a very beautiful fountain in the grounds adorned with allegorical figures. This was purchased by Edward Meldrum, of oil fame, one of the principals of Bathgate Chemical Works, and rebuilt in his grounds at Dechmont Castle.

The Deans Mansion House.

Boghall House, or the Deans Mansion House, the seat of the Norvells, was situated at the Deans, and when taken down the material was used for building the present farm steading. The old mansion was somewhat similar to the one at Kirkton, but on a much larger scale, and consequently much more commodious and in a better state of preservation. It is not known how old the building was, but it dates as far back as 1726, as seen by a tombstone within the church at the Old Churchyard. The family lineage from that date is briefly traced down to 1860, on the death of the last of the family — Mary Margaret Swindell Norvell.

Professor Diney and the Herd Laddie.

Let me add the following additional story about Professor Diney, which was recently related to me by an old Bathgate native, who is now resident in Edinburgh, and was a spectator of the scene some sixty years ago.

Dr Kirk was making his usual daily call at Avonbridge, and was due back. Diney was very unsettled that day, always coming out to the door of his house, shading his eyes and looking towards the Lower Station. At last when the Doctor did arrive, Diney hailed him thus — “Great news the day, Doctor; great news the day!” “Oh, what is the news the day, Professor?” the Doctor asked. Diney replied — “I have had a letter from James Wyllie, the Herd Laddie, who is going to pay a visit to our town, and he is coming to my house. We will have to invite all the noted draughts players in the town to meet him.”

It was a great day for both Diney and Sandy Christie when the Herd Laddie arrived. The house went like “a running fair” the whole day, all the noted draughts players coming to have a game with the champion, Sandy Christie “running the cutter” to keep up the enthusiasm. He had not far to run, his licensed grocer’s shop being next door.

It is not recorded how much the Herd Laddie made on the visit, but he sold a number of his books at 3d each, giving a description of his life.

“Raisin’ the Wind.”

Before concluding let me give you an illustration of a Bathgate worthy’s method of “raising the wind.” This character was addicted to “mountain dew.” He was a schoolmate of Sir James Young Simpson, who, on his visits to Bathgate, used to talk to him about his school-fellow. “Willie” got a good education, and was an excellent Latin scholar. He was well versed with the poets, whom he quoted when appealing for aid. The “Aleck” referred to in the verses was Aleck Hamilton of Kirkroads. This is how he proceeded —

“Not a drum was heard!” Oh, what will I say?

Wad ye len’ me the length o’ a tipence? If ever ye saved a puir mortal man’s life,

O try, man, and mak’ it a thripence. Ye see hoo I’m trumlin’ wi palsy and fricht;

Man a penny’s for nae use ava’ — It’ll only buy Finish, Coffin-Montin’, or Hard,

Or yer All-Sorts, or Speel up the Wa’. Burns is jist Burns! Though he’s clever eneuch,

Yet his poems they never can cope Wi’ Cawmil’s Last Man, or his Brave Soldier’s Dream,

Or Lochiel, or The Pleasures o’ Hope. I ne’er liket Byron, for a’ his fine words,

He’s sae fu’ o’ Tom Payne and Voltaire, And aye findin’ faut wi’ a’ things but himsel’

In Childe Harold, Juan, and Corsair. Hae mercy on me, for I’m likin’ tae fent!

Wad ye mak’ me some gruel or tea? And Aleck, oh Aleck! wad ye fin’ a’ yer pouch,

And gie me anither bawbee.